A rectangular box at the top to have volume of 32cubic feet. Find the dimensions of the box requiring least material for it"s construction


1
Expert's answer
2021-02-03T00:00:57-0500

Step-by-step explanation:


Volume = LBH = 32 cubic feet

To have minimum surface area L = B

"\rightarrow" Volume = L²H = 32

"\rightarrow" H = "\dfrac{32}{L^2}"

Surface Area A = LB + 2( L + B) H

A = L² + 2(L + L)("\dfrac{32}{L^2}")

A= L² + "\dfrac{128}{L}"

"\dfrac{dA}{dL}=2L-\dfrac{128}{L^2}"

"2L-\dfrac{128}{L^2}=0"


"\rightarrow" L³ = 64

"\rightarrow" L = 4


"\dfrac{d^2A}{dL^2}=2+\dfrac{256}{L^3}" ( + Ve)


Hence Area is minimum at L = 4


Surface Area = "L^2+\dfrac{128}{L^2}=4^2+\dfrac{128}{4}=16+32=48\;feet^2"

H = "\dfrac{32}{L^2}=\dfrac{32}{4^2}=2"


Dimension of Box = 4 * 4 * 2 feet


Answer:

Thus, 48 square feet is the least amount of material for its construction.